OpenStudy (anonymous):

Single Variable Calculus 18.01SC @ 36:41mins in; the professor says you can find the y-intercept when you put it into the equation of a tangent line y-y0=(-1/x^2)(x-x0). Since f(x) = y then f(x0) = 1/x. Thus the eq. of the tangent line becomes y-(1/x) = (-1/x^2)(0-x0) The y-intercept SHOULD come out to 2y0 but I end up getting y=2/x0. Can someone help with this? EDIT: This is from the triangle area problem from the video. I tried to find the area of the triangle and I get 2 when my y=2/x0 which is what the professor got as well. What am I doing wrong?

OpenStudy (anonymous):

The equation of the tangent line is: \[(y - y_{0}) = \left( \frac{ -1 }{x_{0}^{2}} \right) (x - x _{0})\] To find the y-intercept, plug in x = 0 to get: \[(y - y_{0}) = \left( \frac{ -1 }{x_{0}^{2}} \right) (0 - x _{0})\] or \[(y - y_{0}) = \frac{ 1 }{x_{0}} \] Since the point (x0, y0) lies on the curve y = 1/x we know that \[y _{0} = \frac{ 1 }{ x _{0} }\] Substituting into the previous equation we get: \[(y - y _{0}) = y_{0} \] \[y = 2y _{0}\] So the y-intercept is 2y0. This is in fact the same thing as 2/x0 so both will give you the correct area of the triangle.

OpenStudy (anonymous):

LOL that was so obvious and stupid of me. Thanks Zigs!